A wire M of diameter d is joint to another wire N of the same material. The diameter of N is half that of M but the two wires have the same length, If a steady potential difference is applied across the ends of the composite wire, which of the following statements is correct?


A. The current in N is half of the current in M


B. The drift velocity of the electrons in N is four times that of the electrons in M.


C. The number of charge carriers per unit length in M is twice that in N


D. The drift velocity of the electrons in M is twice that of the electrons in N


1
Expert's answer
2021-08-18T11:08:27-0400

"v = \frac{I}{enS}"

"I_N=I_M=I(\text{since the serial connection of the conductors})"

"v_N= \frac{I}{enS_N}"

"v_M= \frac{I}{enS_M}"

"\frac{v_N}{v_M}= \frac{S_M}{S_N};"


"\frac{S_M}{S_N}= \frac{\pi\frac{d_M^2}{4}}{\pi\frac{d_N^2}{4}}=4"

"\frac{v_N}{v_M}=4"


"\text{Answer: B. }"